# On algebraic properties of some non-prime ideals of collections of cells

### v2

@@ 4. Cohen-Macaulay property and Hilbert series for closed path polyominoes having zig-zag walks
Discussion 4.1.
If $\mathcal{P}$ is a closed path polyomino having zig-zag walks, by [3, Proposition 6.1] $\mathcal{P}$ does not contain any $L$-configuration or ladders with at least three steps. So, any two maximal inner intervals of $\mathcal{P}$ intersect themselves in the cells displayed in Figure 14a or 14b (up to reflections or rotations). Furthermore, by [6] we know that there exists a monomial order such that the Gröbner basis is quadratic and squarefree. The monomial order that provides this property can be built following [6, Algorithm 4.1]. Following this construction, we obtain that for all inner intervals containing the vertex $m$ as in Figure 14a, (up to reflection and rotation), the initial term of the related inner 2-minors contain the variable $x_{m}$ in its support. From now onward we denote by $M_{\mathcal{P}}$ the set of all vertices $m$ in $\mathcal{P}$ as in Figure 14a, up to rotations and reflections, and by $C_{m}$ the cell labeled with $X$ with reference to the same figure.
@@
Let $\mathcal{P}$ be a closed path with a zig-zag walk and $\prec$ be the monomial order on $S_{\mathcal{P}}$ defined by [6, Algorithm 4.1], that provides a quadratic Gröbner basis for $I_{\mathcal{P}}$. For all $i\in[r]$, with reference to Set-up 4.2, since $V(\mathcal{Q}_{i})\subset V(\mathcal{P})$ and $V(\mathcal{Q}_{i}^{\prime})\subset V(\mathcal{P})$, the monomial order $\prec$ can be considered also in $S_{\mathcal{Q}_{i}}$ and $S_{\mathcal{Q}^{\prime}_{i}}$. In particular, $\prec$ trivially preserves the same initial terms in the ideals $I_{\mathcal{P}}$, $I_{\mathcal{Q}_{i}}$ and $I_{\mathcal{Q}^{\prime}_{i}}$. Furthermore, by [7, Proposition 4.1], $\prec$ provides a quadratic squarefree Gröbner basis also for $I_{\mathcal{Q}_{i}}$ and $I_{\mathcal{Q}_{i}^{\prime}}$. So, with abuse of notation, we refer to $\mathrm{in}(I_{\mathcal{Q}_{i}})$, $\mathrm{in}(I_{\mathcal{Q}_{i-1}})$ and $\mathrm{in}(I_{\mathcal{Q}^{\prime}_{i}})$ considering the monomial order $\prec$.
@@
Let $\prec$ be the monomial order in $S_{\mathcal{P}}$ defined by [6, Algorithm 4.1], that provides a quadratic Gröbner basis for $I_{\mathcal{P}}$. For all $i\in[r]$, we focus on the initial ideals of $I_{\mathcal{Q}_{i}}$ and $I_{\mathcal{Q}_{i}^{\prime}}$ as explained in Remark 4.3. In particular, let $i\in[r]$ and consider the following exact sequence:
@@
Arguing as in the proof of [7, Proposition 4.3], we can prove that:
@@
So by Propositions 1.5 and 1.7 it follows that for all $i\in\{1\,\ldots,r\}$ it is verified:
@@
and also, by Propositions 1.6, 1.9 and 1.10, we obtain $\mathrm{depth}(K[\mathcal{Q}_{i-1}])\geq\min\{\mathrm{depth}(K[\mathcal{Q}_{i}]),\mathrm{depth}(K[\mathcal{Q}_{i}^{\prime}])+1\}$. In particular, we obtain the first equality of $\mathrm{HP}_{K[\mathcal{P}]}(t)$.
@@
If $\mathcal{P}$ is a non-prime closed path polyomino not containing the configuration in Figure 14b then $\mathcal{Q}_{r}$ is a zig-zag collection and by Theorem 3.6 we have that $K[\mathcal{Q}_{r}]$ is Cohen-Macaulay of dimension $|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|=|V(\mathcal{P})|-|\mathcal{P}|$. Hence $\mathrm{depth}(K[\mathcal{P}])\geq|V(\mathcal{P})|-|\mathcal{P}|$ but we also know that $\mathrm{dim}(K[\mathcal{P}])=|V(\mathcal{P})|-|\mathcal{P}|$ (see [10, Theorem 3.4]), so $K[\mathcal{P}]$ is Cohen-Macaulay of dimension $|V(\mathcal{P})|-|\mathcal{P}|$. ∎
@@
Let $\mathcal{Q}_{r}$ be a collection of cells obtained as in Set-up 4.2. Recall that $\mathcal{Q}_{r}$ is obtained from a closed path $\mathcal{P}$ having zig-zag walks, by removing all cells $X$ in each sub-configurations of $\mathcal{P}$ as in Figure 14a (up to reflections and rotations). Observe that, since $\mathcal{P}$ has not $L$-configurations and ladders of at least three steps, then $\mathcal{Q}_{r}$ can be built as a union of opportune reflections and rotations of the configurations in Figure 16. That is, we can express $\mathcal{Q}_{r}=\bigcup_{k=1}^{q}\mathcal{C}_{k}$, where $\mathcal{C}_{k}$ is a configuration as in Figure 16a or Figure 16b or Figure 16c (up to rotations and reflections), for all $k\in\{1,\ldots,q\}$, for some positive integer $q$. In particular, with reference to the highlighted points in the figures, we build $\mathcal{Q}_{r}$ by gluing those configurations with this procedure: setting the first configuration $\mathcal{C}_{k}$, the second one, say $\mathcal{C}_{l}$ with $l=k+1$, is glued considering an opportune reflection or rotation of it, in such a way that the vertices $c_{k},d_{k}$ and $e_{k}$ in orange of the first configuration are overlapped, respectively, to the vertices $c_{l-1},d_{l-1}$ and $e_{l-1}$ in blue of the second configuration. See Figure 17 for an example of this construction.
@@
In what follows, we often refer to the framework explained in Set-up 4.8, and we made extensive use of the notations introduced there, also with the help of figures.
@@
Denote $s=\max J_{\mathcal{Q}_{r}}$. Observe that $\mathcal{R}_{n_{s},s}$ is a zig-zag collection, so we denote by $I_{1},\ldots,I_{p}$ the set of maximal intervals of such a zig-zag collection, for some integer $p$. We can assume, without loss of generality, that $I_{1}$ is related to the cell interval $[E_{1},B_{1,1}]\setminus\{B_{1,1}\}$, with reference to the notations explained in Set-up 4.8 and to Figures 16a and 16b, assuming also that $I_{1}$ have the same orientation of these figures.
@@
Set $\prec_{\mathcal{Q}_{r}}$ the lexicographic order on $S_{\mathcal{Q}_{r}}$ induced by the total order $<_{\mathcal{Q}_{r}}$. Trivially, the monomial order $\prec_{\mathcal{Q}_{r}}$ satisfies claim (2). We want to prove also that claim (1) holds. Firstly, we prove it for the ideal $I_{\mathcal{Q}_{r}}$.
Let $f,g$ be inner 2-minors of $I_{\mathcal{Q}_{r}}$ and denote by $S(f,g)$ the $S$-polynomial of $f$ and $g$. It is not difficult to see that if $f,g\in I_{\mathcal{R}_{n_{s},s}}$, since $S(f,g)$ reduces to 00 in $I_{\mathcal{R}_{n_{s},s}}$, then $S(f,g)$ reduces to 00 also in $I_{\mathcal{Q}_{r}}$. So, suppose $f\notin I_{\mathcal{R}_{n_{s},s}}$. Hence the inner interval related to $f$ contains the cell $B_{i,k}$ for some $k\in J_{\mathcal{Q}_{r}}$ and $i\in[n_{k}]$. Then, referring to Figure 19, up to reflection an rotations $f$ is one of the following:
@@
    $f=f_{j,i,k}$ with $j\in[\ell]$ and $g=\pm(x_{v_{j_{1}}}x_{h_{j_{2}}}-x_{v_{j_{2}}}x_{h_{j_{1}}})$ with $j_{1},j_{2}\in[\ell]$ or $g=f_{l,i,k}$ with $l\in[\ell]\setminus\{j\}$. In this case $S(f,g)$ reduces to 00 by [6, Remark 1].
@@
    $f=f_{1,i,k}$ and $g=f_{b,i,k}$. Also in this case $S(f,g)$ reduces to 00 by [6, Remark 1].
@@
    $f=f_{j,i,k}$ with $j\in[\ell]\setminus\{1\}$ and $g=f_{b,i,k}$. In this case, by [6, Lemma 2 (condition 1)], we obtain that $S(f,g)$ reduces to 00 if both the following properties holds:
@@
    $f=f_{1,i,k}$ and $g=\pm(x_{h_{1}}x_{u}-x_{a_{i,k}^{(1)}}x_{w})$ with $u,w\in V([A_{i,k},B_{i+1,k}]$ if $i\neq n_{k}$, or $u,w\in V([A_{i,k},D_{k}]$ if $i=n_{k}$, and $\operatorname{in}(g)=x_{h_{1}}x_{u}$. In this case, recall that $x_{z}\prec_{\mathcal{Q}_{r}}x_{b_{i,k}^{(2)}}$ for all $z\notin Y$. Hence, if $u\notin Y$ or $u\in Y$ with $x_{u}\prec_{\mathcal{Q}_{r}}x_{b_{i,k}^{(2)}}$ then $S(f,g)$ reduces to 00 by [6, Lemma 4 (condition 3)], while if $u\in Y$ and $x_{b_{i,k}^{(2)}}\prec_{\mathcal{Q}_{r}}x_{u}$ then $S(f,g)$ reduces to 00 by [6, Lemma 4 (condition 1)].
@@
Observe that, since by assumptions $I_{\mathcal{P}}$ has a quadratic Gröbner basis with respect to $\prec$, $x_{b_{r}}$ does not divide any generator of $\operatorname{in}(I_{\mathcal{P}})$ different by $x_{a_{i}}x_{b_{r}}$, for $i\in[r-1]$. In fact, assume $f=x_{b_{r}}x_{v}-x_{z}x_{w}\in I_{\mathcal{P}}$ for some $v,w,z\in V(\mathcal{P})$ such that $v\in V(\mathcal{P})\setminus\{a,a_{1},\ldots,a_{r}\}$ and $\operatorname{in}(f)=x_{b_{r}}x_{v}$. So, considering $g=x_{a}x_{b_{r}}-x_{b}x_{a_{r}}\in I_{\mathcal{P}}$, we have $\gcd(\operatorname{in}(f),\operatorname{in}(g))\neq 0$. Without loss of generality, assume $z$ is in the same horizontal edge interval of $b_{r}$, while $w$ is in the same vertical edge interval of $b_{r}$. As consequence, assuming the $S$-polynomial of $f$ and $g$ reduces to zero, by [6, Lemma 6] we have $[z,a_{r}]$ or $[w,a]$ is an inner interval, a contradiction.
@@
Hence, by Proposition 1.8, we have $\mathrm{HP}_{S_{\mathcal{R}}/(\operatorname{in}(I_{\mathcal{R}}),x_{b_{r}})}(t)=(1-t)\cdot\mathrm{HP}_{S_{\mathcal{R}}/\operatorname{in}(I_{\mathcal{R}})}(t)$.
Therefore, by Proposition 1.5 and Proposition 1.7 we obtain $\mathrm{HP}_{K[\mathcal{P}]}(t)=\mathrm{HP}_{K[\mathcal{R}]}(t)+\frac{t}{(1-t)^{r}}\mathrm{HP}_{K[\mathcal{R}]}(t)$, that is the first claim.
For the second claim, by Proposition 1.10 we have $\operatorname{depth}(S_{\mathcal{P}}/\operatorname{in}(I_{\mathcal{P}}))\geq\{\operatorname{depth}(S_{\mathcal{Q}}/\operatorname{in}(I_{\mathcal{Q}})),\operatorname{depth}(S_{\mathcal{R}}/(\operatorname{in}(I_{\mathcal{R}}),x_{b_{r}})\otimes K[x_{a},x_{a_{1}},\ldots,x_{a_{r}}])\}$. Hence, by Propositions 1.6 and 1.9 we obtain $\operatorname{depth}(K[\mathcal{P}])\geq\min\{\operatorname{depth}(K[\mathcal{Q}]),\operatorname{depth}(K[\mathcal{R}])+r\}$.
@@
Observe that, up to rotations and reflections, the configuration of cells given by $\mathcal{E}_{i,k}\cup\{A_{i,k}\}$ determine a sub-configuration of $\mathcal{S}_{i,k}$ as in Figure 20. Moreover, observe that $\mathcal{S}_{i,k}\setminus\{B_{i,k}\}=\mathcal{R}_{i,k}$ and $\mathcal{S}_{i,k}\setminus(\mathcal{E}_{i,k}\cup\{A_{i,k}\})=\mathcal{R}_{i,k}^{\prime}$. Furthermore, by Lemma 4.10, we have that $\mathcal{S}_{i,k}$ satisfies all conditions of Lemma 4.11 and $\mathcal{R}_{i,k}^{\prime}$ is a simple collection of cells. Therefore $\mathrm{HP}_{K[\mathcal{S}_{i,k}]}(t)=\mathrm{HP}_{K[\mathcal{R}_{i,k}]}(t)+\frac{t\cdot h_{K[\mathcal{R}^{\prime}_{i,k}]}(t)}{(1-t)^{|V(\mathcal{S}_{i,k})|-|\mathcal{S}_{i,k}|}}$. and $\operatorname{depth}(K[\mathcal{S}_{i,k}])\geq\min\{\operatorname{depth}(K[\mathcal{R}_{i,k}]),|V(\mathcal{S}_{i,k})|-|\mathcal{S}_{i,k}|\}$.
So, denoting $J_{\mathcal{Q}_{r}}=\{k_{1},\ldots,k_{s}\}$, starting from $\mathcal{Q}_{r}$ we obtain:
@@
Observing that $|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|=|V(\mathcal{R}_{i,k})|-|\mathcal{R}_{i,k}|$ for all $k\in J_{\mathcal{Q}_{r}}$ and for all $i\in[n_{k}]$ (in fact $\mathcal{S}_{i,k}$ differs from $\mathcal{R}_{i,k}$ only for a vertex and cells), we obtain $\operatorname{depth}(K[\mathcal{Q}_{r}])\geq\min\{\operatorname{depth}(K[\mathcal{R}_{n_{k_{s}},k_{s}}]),|V(Q_{r})|-|\mathcal{Q}_{r}|\}$. Finally, observe that $\mathcal{R}_{n_{k_{s}},k_{s}}$ is a zig-zag collection. Hence, by Theorem 3.6, we have that $K[\mathcal{R}_{n_{k_{s}},k_{s}}]$ is a Cohen-Macaulay domain of dimension $|V(\mathcal{R}_{n_{k_{s}},k_{s}})|-|\mathcal{R}_{n_{k_{s}},k_{s}}|=|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|$, obtaining $\operatorname{depth}(K[\mathcal{Q}_{r}])\geq|V(Q_{r})|-|\mathcal{Q}_{r}|$
@@
By Lemma 4.5 we have $\mathrm{depth}(K[\mathcal{P}])\geq\min\{\mathrm{depth}(K[\mathcal{Q}_{r}]),|V(\mathcal{P})|-|\mathcal{P}|\}$ and by Lemma 4.12 $\mathrm{depth}(K[\mathcal{Q}_{r}])\geq|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|$. Moreover, by the proof of Lemma 4.5 we have also $|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|=|V(\mathcal{P})|-|\mathcal{P}|$. Hence, $\operatorname{depth}(K[\mathcal{P}])\geq|V(\mathcal{P})|-|\mathcal{P}|$. By [10, Theorem 3.4] we also know that $\dim(K[\mathcal{P}])=|V(\mathcal{P})|-|\mathcal{P}|$, so $K[\mathcal{P}]$ is Cohen-Macaulay with the given dimension. ∎

### v3

@@ 4. Cohen-Macaulay property and Hilbert-Poincaré series of closed paths with zig-zag walks
Furthermore, by [7] we know that there exists a monomial order such that the Gröbner basis is quadratic and squarefree. The monomial order that provides this property can be built following [7, Algorithm 4.1]. Following this construction, we obtain that for all inner intervals containing the vertex $m$ as in Figure 10a, (up to reflection and rotation), the initial term of the related inner 2-minors contain the variable $x_{m}$ in its support. From now onward, we denote by $M_{\mathcal{P}}$ the set of all vertices $m$ in $\mathcal{P}$ as in Figure 10a, up to rotations and reflections, and by $C_{m}$ the cell labeled with $X$ with reference to the same figure.
@@
Let $\mathcal{P}$ be a closed path with a zig-zag walk and $\prec$ be the monomial order on $S_{\mathcal{P}}$ defined by [7, Algorithm 4.1], that provides a quadratic Gröbner basis for $I_{\mathcal{P}}$. For all $i\in[r]$, with reference to Set-up 4.2, since $V(\mathcal{Q}_{i})\subset V(\mathcal{P})$ and $V(\mathcal{Q}_{i}^{\prime})\subset V(\mathcal{P})$, the monomial order $\prec$ can be considered also in $S_{\mathcal{Q}_{i}}$ and $S_{\mathcal{Q}^{\prime}_{i}}$. In particular, $\prec$ trivially preserves the same initial terms in the ideals $I_{\mathcal{P}}$, $I_{\mathcal{Q}_{i}}$ and $I_{\mathcal{Q}^{\prime}_{i}}$. Furthermore, by [8, Proposition 4.1], $\prec$ provides a quadratic squarefree Gröbner basis also for $I_{\mathcal{Q}_{i}}$ and $I_{\mathcal{Q}_{i}^{\prime}}$. So, with abuse of notation, we refer to $\mathrm{in}(I_{\mathcal{Q}_{i}})$, $\mathrm{in}(I_{\mathcal{Q}_{i-1}})$ and $\mathrm{in}(I_{\mathcal{Q}^{\prime}_{i}})$ considering the monomial order $\prec$.
@@
Let $\prec$ be the monomial order in $S_{\mathcal{P}}$ defined by [7, Algorithm 4.1], that provides a quadratic Gröbner basis for $I_{\mathcal{P}}$. For all $i\in[r]$, we focus on the initial ideals of $I_{\mathcal{Q}_{i}}$ and $I_{\mathcal{Q}_{i}^{\prime}}$ as explained in Remark 4.3. In particular, let $i\in[r]$ and consider the following exact sequence:
@@
Arguing as in the proof of [8, Proposition 4.3], we can prove that:
@@
So by [37, Corollary 3.3.15] and [37, Lemmas 5.1.2 and 5.1.3] it follows that for all $i\in\{1\,\ldots,r\}$ it is verified:
@@
and also, by Propositions 3.6, [37, Lemma 3.1.34] and [3, Proposition 1.2.9], we obtain $\mathrm{depth}(K[\mathcal{Q}_{i-1}])\geq\min\{\mathrm{depth}(K[\mathcal{Q}_{i}]),\mathrm{depth}(K[\mathcal{Q}_{i}^{\prime}])+1\}$. In particular, we obtain the first equality of $\mathrm{HP}_{K[\mathcal{P}]}(t)$.
@@
If $\mathcal{P}$ is a non-prime closed path polyomino not containing the configuration in Figure 10b then $\mathcal{Q}_{r}$ is a zig-zag collection and by Theorem 3.7 we have that $K[\mathcal{Q}_{r}]$ is Cohen-Macaulay of dimension $|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|=|V(\mathcal{P})|-|\mathcal{P}|$. Hence $\mathrm{depth}(K[\mathcal{P}])\geq|V(\mathcal{P})|-|\mathcal{P}|$ but we also know that $\mathrm{dim}(K[\mathcal{P}])=|V(\mathcal{P})|-|\mathcal{P}|$ (see [11, Theorem 3.4]), so $K[\mathcal{P}]$ is Cohen-Macaulay of dimension $|V(\mathcal{P})|-|\mathcal{P}|$. ∎
@@
Let $\mathcal{Q}_{r}$ be a collection of cells obtained as in Set-up 4.2. Recall that $\mathcal{Q}_{r}$ is obtained from a closed path $\mathcal{P}$ having zig-zag walks, by removing all cells $X$ in each sub-configurations of $\mathcal{P}$ as in Figure 10a (up to reflections and rotations). Observe that, since $\mathcal{P}$ has not $L$-configurations and ladders of at least three steps, then $\mathcal{Q}_{r}$ can be built as a union of opportune reflections and rotations of the configurations in Figure 12. That is, we can express $\mathcal{Q}_{r}=\bigcup_{k=1}^{q}\mathcal{C}_{k}$, where $\mathcal{C}_{k}$ is a configuration as in Figure 12a or Figure 12b or Figure 12c (up to rotations and reflections), for all $k\in\{1,\ldots,q\}$, for some positive integer $q$. In particular, with reference to the points highlighted in the figures, we build $\mathcal{Q}_{r}$ by gluing those configurations with this procedure: setting the first configuration $\mathcal{C}_{k}$, the second one, say $\mathcal{C}_{l}$ with $l=k+1$, is glued considering an opportune reflection or rotation of it, in such a way that the vertices $c_{k},d_{k}$ and $e_{k}$ in orange of the first configuration are overlapped, respectively, to the vertices $c_{l-1},d_{l-1}$ and $e_{l-1}$ in blue of the second configuration. See Figure 13 for an example of this construction.
@@
In what follows, we often refer to the framework explained in Set-up 4.8, and we made extensive use of the notation introduced there, also with the help of figures.
@@
Denote $s=\max J_{\mathcal{Q}_{r}}$. Observe that $\mathcal{R}_{n_{s},s}$ is a zig-zag collection, so we denote by $I_{1},\ldots,I_{p}$ the set of maximal intervals of such a zig-zag collection, for some integer $p$. We can assume, without loss of generality, that $I_{1}$ is related to the cell interval $[E_{1},B_{1,1}]\setminus\{B_{1,1}\}$, with reference to the notations explained in Set-up 4.8 and to Figures 12a and 12b, assuming also that $I_{1}$ have the same orientation of these figures.
@@
Set $\prec_{\mathcal{Q}_{r}}$ the lexicographic order on $S_{\mathcal{Q}_{r}}$ induced by the total order $<_{\mathcal{Q}_{r}}$. Trivially, the monomial order $\prec_{\mathcal{Q}_{r}}$ satisfies claim (2). We also want to prove that claim (1) holds. Firstly, we prove it for the ideal $I_{\mathcal{Q}_{r}}$.
Let $f,g$ be inner 2-minors of $I_{\mathcal{Q}_{r}}$ and denote by $S(f,g)$ the $S$-polynomial of $f$ and $g$. It is not difficult to see that if $f,g\in I_{\mathcal{R}_{n_{s},s}}$, since $S(f,g)$ reduces to 00 in $I_{\mathcal{R}_{n_{s},s}}$, then $S(f,g)$ reduces to 00 also in $I_{\mathcal{Q}_{r}}$. So, suppose $f\notin I_{\mathcal{R}_{n_{s},s}}$. Hence the inner interval related to $f$ contains the cell $B_{i,k}$ for some $k\in J_{\mathcal{Q}_{r}}$ and $i\in[n_{k}]$. Then, referring to Figure 15, up to reflection and rotations, $f$ is one of the following:
@@
    $f=f_{j,i,k}$ with $j\in[\ell]$ and $g=\pm(x_{v_{j_{1}}}x_{h_{j_{2}}}-x_{v_{j_{2}}}x_{h_{j_{1}}})$ with $j_{1},j_{2}\in[\ell]$ or $g=f_{l,i,k}$ with $l\in[\ell]\setminus\{j\}$. In this case $S(f,g)$ reduces to 00 by [7, Remark 1].
@@
    $f=f_{1,i,k}$ and $g=f_{b,i,k}$. In this case also, $S(f,g)$ reduces to 00 by [7, Remark 1].
@@
    $f=f_{j,i,k}$ with $j\in[\ell]\setminus\{1\}$ and $g=f_{b,i,k}$. In this case, by [7, Lemma 2 (condition 1)], we obtain that $S(f,g)$ reduces to 00 if both the following properties holds:
@@
    $f=f_{1,i,k}$ and $g=\pm(x_{h_{1}}x_{u}-x_{a_{i,k}^{(1)}}x_{w})$ with $u,w\in V([A_{i,k},B_{i+1,k}]$ if $i\neq n_{k}$, or $u,w\in V([A_{i,k},D_{k}]$ if $i=n_{k}$, and $\operatorname{in}(g)=x_{h_{1}}x_{u}$. In this case, recall that $x_{z}\prec_{\mathcal{Q}_{r}}x_{b_{i,k}^{(2)}}$ for all $z\notin Y$. Hence, if $u\notin Y$ or $u\in Y$ with $x_{u}\prec_{\mathcal{Q}_{r}}x_{b_{i,k}^{(2)}}$ then $S(f,g)$ reduces to 00 by [7, Lemma 4 (condition 3)], while if $u\in Y$ and $x_{b_{i,k}^{(2)}}\prec_{\mathcal{Q}_{r}}x_{u}$ then $S(f,g)$ reduces to 00 by [7, Lemma 4 (condition 1)].
@@
Observe that, since by assumptions $I_{\mathcal{P}}$ has a quadratic Gröbner basis with respect to $\prec$, $x_{b_{r}}$ does not divide any generator of $\operatorname{in}(I_{\mathcal{P}})$ different by $x_{a_{i}}x_{b_{r}}$, for $i\in[r-1]$. In fact, assume $f=x_{b_{r}}x_{v}-x_{z}x_{w}\in I_{\mathcal{P}}$ for some $v,w,z\in V(\mathcal{P})$ such that $v\in V(\mathcal{P})\setminus\{a,a_{1},\ldots,a_{r}\}$ and $\operatorname{in}(f)=x_{b_{r}}x_{v}$. So, considering $g=x_{a}x_{b_{r}}-x_{b}x_{a_{r}}\in I_{\mathcal{P}}$, we have $\gcd(\operatorname{in}(f),\operatorname{in}(g))\neq 0$. Without loss of generality, assume $z$ is in the same horizontal edge interval of $b_{r}$, while $w$ is in the same vertical edge interval of $b_{r}$. As a consequence, assuming that the $S$-polynomial of $f$ and $g$ reduces to zero, by [7, Lemma 6] we have $[z,a_{r}]$ or $[w,a]$ is an inner interval, a contradiction.
@@
Hence, by [37, Lemma 5.1.11], we have $\mathrm{HP}_{S_{\mathcal{R}}/(\operatorname{in}(I_{\mathcal{R}}),x_{b_{r}})}(t)=(1-t)\cdot\mathrm{HP}_{S_{\mathcal{R}}/\operatorname{in}(I_{\mathcal{R}})}(t)$. Therefore, by [37, Corollary 3.3.15] and [37, Lemmas 5.1.2 and 5.1.3] we obtain $\mathrm{HP}_{K[\mathcal{P}]}(t)=\mathrm{HP}_{K[\mathcal{R}]}(t)+\frac{t}{(1-t)^{r}}\mathrm{HP}_{K[\mathcal{R}]}(t)$, that is the first claim.
For the second claim, by [3, Proposition 1.2.9] we have $\operatorname{depth}(S_{\mathcal{P}}/\operatorname{in}(I_{\mathcal{P}}))\geq\{\operatorname{depth}(S_{\mathcal{Q}}/\operatorname{in}(I_{\mathcal{Q}})),\operatorname{depth}(S_{\mathcal{R}}/(\operatorname{in}(I_{\mathcal{R}}),x_{b_{r}})\otimes K[x_{a},x_{a_{1}},\ldots,x_{a_{r}}])\}$. Hence, by Propositions 3.6 and [37, Lemma 3.1.34] we obtain $\operatorname{depth}(K[\mathcal{P}])\geq\min\{\operatorname{depth}(K[\mathcal{Q}]),\operatorname{depth}(K[\mathcal{R}])+r\}$.
@@
Observe that, up to rotations and reflections, the configuration of cells given by $\mathcal{E}_{i,k}\cup\{A_{i,k}\}$ determine a sub-configuration of $\mathcal{S}_{i,k}$ as in Figure 16. Moreover, observe that $\mathcal{S}_{i,k}\setminus\{B_{i,k}\}=\mathcal{R}_{i,k}$ and $\mathcal{S}_{i,k}\setminus(\mathcal{E}_{i,k}\cup\{A_{i,k}\})=\mathcal{R}_{i,k}^{\prime}$. Furthermore, by Lemma 4.10, we have that $\mathcal{S}_{i,k}$ satisfies all conditions of Lemma 4.11 and $\mathcal{R}_{i,k}^{\prime}$ is a simple collection of cells. Therefore $\mathrm{HP}_{K[\mathcal{S}_{i,k}]}(t)=\mathrm{HP}_{K[\mathcal{R}_{i,k}]}(t)+\frac{t\cdot h_{K[\mathcal{R}^{\prime}_{i,k}]}(t)}{(1-t)^{|V(\mathcal{S}_{i,k})|-|\mathcal{S}_{i,k}|}}$. and $\operatorname{depth}(K[\mathcal{S}_{i,k}])\geq\min\{\operatorname{depth}(K[\mathcal{R}_{i,k}]),|V(\mathcal{S}_{i,k})|-|\mathcal{S}_{i,k}|\}$. So, denote $J_{\mathcal{Q}_{r}}=\{k_{1},\ldots,k_{s}\}$. Starting from $\mathcal{Q}_{r}$ we obtain:
@@
Observing that $|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|=|V(\mathcal{R}_{i,k})|-|\mathcal{R}_{i,k}|$ for all $k\in J_{\mathcal{Q}_{r}}$ and for all $i\in[n_{k}]$ (in fact, $\mathcal{S}_{i,k}$ differs from $\mathcal{R}_{i,k}$ only for a vertex and cells), we obtain $\operatorname{depth}(K[\mathcal{Q}_{r}])\geq\min\{\operatorname{depth}(K[\mathcal{R}_{n_{k_{s}},k_{s}}]),|V(Q_{r})|-|\mathcal{Q}_{r}|\}$. Finally, observe that $\mathcal{R}_{n_{k_{s}},k_{s}}$ is a zig-zag collection. Hence, by Theorem 3.7, we have that $K[\mathcal{R}_{n_{k_{s}},k_{s}}]$ is a Cohen-Macaulay domain of dimension $|V(\mathcal{R}_{n_{k_{s}},k_{s}})|-|\mathcal{R}_{n_{k_{s}},k_{s}}|=|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|$, obtaining $\operatorname{depth}(K[\mathcal{Q}_{r}])\geq|V(Q_{r})|-|\mathcal{Q}_{r}|$
@@
By Lemma 4.5 we have $\mathrm{depth}(K[\mathcal{P}])\geq\min\{\mathrm{depth}(K[\mathcal{Q}_{r}]),|V(\mathcal{P})|-|\mathcal{P}|\}$ and by Lemma 4.12 $\mathrm{depth}(K[\mathcal{Q}_{r}])\geq|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|$. Moreover, by the proof of Lemma 4.5 we also have $|V(\mathcal{Q}_{r})|-|\mathcal{Q}_{r}|=|V(\mathcal{P})|-|\mathcal{P}|$. Hence, $\operatorname{depth}(K[\mathcal{P}])\geq|V(\mathcal{P})|-|\mathcal{P}|$. By [11, Theorem 3.4] we also know that $\dim(K[\mathcal{P}])=|V(\mathcal{P})|-|\mathcal{P}|$, so $K[\mathcal{P}]$ is Cohen-Macaulay with the given dimension. ∎